three tea bags are needed to make a gallon of iced tea. How many tea bags are needed to make four gallons

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

3*4=12 tea bags

Answer by stanbon(75887)

three tea bags are needed to make a gallon of iced tea. How many tea bags are needed to make four gallons
-----------------
Proportion:
x/4 = 3/1
---
Cross-multiply to solve for "x":
x = 12 tea bags needed to make 4 gallons
============================================
Cheers,
Stan H.